# Brain-Inspired Stepwise Patch Merging for Vision Transformers

### v1

@@ Abstract
The hierarchical architecture has become a mainstream design paradigm for Vision Transformers (ViTs), with Patch Merging serving as the pivotal component that transforms a columnar architecture into a hierarchical one. Drawing inspiration from the brain’s ability to integrate global and local information for comprehensive visual understanding, we propose a novel technique called Stepwise Patch Merging (SPM), which enhances the subsequent attention mechanism’s ability to ’see’ better. SPM comprises two critical modules: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE). The MSA module integrates multi-scale features to enrich feature representation, while the GLE module focuses on refining local detail extraction, thus achieving an optimal balance between long-range dependency modeling and local feature enhancement. Extensive experiments conducted on benchmark datasets, including ImageNet-1K, COCO, and ADE20K, demonstrate that SPM significantly improves the performance of various models, particularly in dense prediction tasks such as object detection and semantic segmentation. These results underscore the efficacy of SPM in enhancing model accuracy and robustness across a wide range of computer vision tasks.
Introduction
@@
Transformers have demonstrated remarkable advancements in natural language processing (NLP) (Vaswani et al. 2017; Devlin et al. 2018), and their application has recently extended significantly into the computer vision (CV) domain (Dosovitskiy et al. 2020; Touvron et al. 2021). To enhance their adaptability to downstream tasks, hierarchical vision transformers (HVTs) (Wang et al. 2021; Liu et al. 2021; Wu et al. 2021; Yan et al. 2021) have been developed. These architectures draw inspiration from the pyramid structure utilized in convolutional neural networks (CNNs) (Krizhevsky, Sutskever, and Hinton 2012; He et al. 2016; Tan and Le 2019; Lin et al. 2017). In HVTs, transformer blocks are segmented into multiple stages, resulting in a progressive reduction of feature map sizes and an increase in the number of channels as the network depth increases.
@@
Figure 1: Overview of the proposed Stepwise Patch Merging (SPM) framework built upon the hierarchical vision transformers. The SPM framework comprises two sequential modules: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE).
@@
HVTs commonly utilize either standard convolutional layers or linear projection layers to amalgamate adjacent tokens (Wang et al. 2021; Liu et al. 2021; Ren et al. 2022), with the objective of generating hierarchical feature maps. Nevertheless, fixed-grid methods can limit the representational capacity of vision transformers in modeling geometric transformations, as not every pixel contributes equally to an output unit (Luo et al. 2016). To overcome this limitation, adaptive methods have been proposed to derive more informative downsampled tokens for subsequent processing. For example, LIT (Pan et al. 2022), drawing inspiration from deformable convolutions (Dai et al. 2017; Zhu et al. 2019), learns a grid of offsets to adaptively adjust spatial sampling locations for merging neighboring patches from a sub-window in a feature map. Similarly, TCformer (Zeng et al. 2022) employs a variant of the k-nearest neighbor-based density peaks clustering algorithm (DPC-KNN) (Du, Ding, and Jia 2016) to aggregate redundant patches, generating more patches on the target object to capture additional information. HAFA (Chen et al. 2023) integrates the methodologies of LIT and TCformer, predicting offsets to adjust the sampling center of patches in the shallow layers, while employing clustering in the deeper layers to group patches with similar semantics in the feature space. However, these methods face several common challenges. They often exhibit limited capacity for modeling long-distance relationships and suffer from a loss of spatial information due to the clustering process. Additionally, the clustering algorithms used are typically not amenable to end-to-end training, leading to inefficiencies. The integration of multiple modules, as seen in HAFA, further complicates their generalizability across different applications.
@@
In the brain, the primary visual cortex (V1), integral to initial visual processing, houses neurons with relatively small receptive fields that are crucial for detecting fine, localized visual features such as edges and orientations. As visual information propagates to higher cortical areas like V2, V3, and V4, neurons with increasingly larger receptive fields integrate these initial perceptions, facilitating the recognition of more complex patterns and broader contextual elements (Hubel and Wiesel 1962; Livingstone and Hubel 1988; Zeki 1978). Additionally, the visual cortex benefits from a dynamic feedback system where higher-order areas like the inferotemporal cortex (IT) provide contextual modulation to lower areas. This top-down modulation is essential for refining the perception of local features within their broader environmental matrix, enhancing both the accuracy and relevance of visual processing (Gilbert and Li 2013; Bullier 2001; Lamme and Roelfsema 2000).
@@
Inspired by the nuanced neurobiological mechanisms of the human visual cortex, particularly the orchestrated activities across various cortical areas, we introduce Stepwise Patch Merging (SPM), a novel approach designed to enhance the receptive field while preserving local details. SPM framework consists of two sequential stages: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE). In the MSA stage, spatial dimensions are preserved while channel dimensions are increased to a designated size. This process aggregates multi-scale information, enriching the semantic content to accommodate the increased capacity of the feature map. Subsequently, the GLE stage reduces the spatial dimensions of the feature map while maintaining the channel dimensions. Given that the input to GLE already contains rich semantic information, this stage emphasizes local information, optimizing it for downstream dense prediction tasks such as object detection and semantic segmentation. The distinct focus and reasonable division of labor between the MSA and GLE modules ensure that the SPM architecture serves as a flexible, drop-in replacement for existing hierarchical vision transformers.
@@
  • We propose an innovative technique termed Stepwise Patch Merging (SPM), which serves as a plug-in replacement within hierarchical vision transformer architectures, leading to substantial performance enhancements.
@@
Related Work
Vision Transformer
The Vision Transformer (ViT) (Dosovitskiy et al. 2020) revolutionized visual tasks by introducing the transformer architecture to computer vision. ViT segments images into non-overlapping patches, projects these patches linearly into token sequences, and processes them using a transformer encoder. ViT models have demonstrated superior performance in image classification and other downstream tasks, surpassing CNNs (He et al. 2016; Krizhevsky, Sutskever, and Hinton 2012; LeCun, Bengio et al. 1995) when trained with large-scale pretraining datasets and advanced training methodologies. Motivated by the success of CNNs and the necessity to address dense prediction tasks, researchers have incorporated the feature pyramid structure within transformers. This innovation has led to the development and widespread adoption of HVTs (Dong et al. 2022; Ren et al. 2022; Fan et al. 2023; Guo et al. 2022; Hou et al. 2022).
Hierarchical Feature Representation
Hierarchical feature representation plays a pivotal role in dense prediction tasks, prompting extensive research in this domain. Existing approaches can be broadly categorized into fixed-grid and dynamic feature-based methods. Fixed-grid methods, exemplified by works such as PVT (Wang et al. 2021) and Swin (Liu et al. 2021), merge patches within adjacent windows using 2D convolution. In contrast, dynamic methods, such as DynamicViT (Rao et al. 2021) adaptively extract features by eliminating redundant patches and retaining essential ones, thereby forming hierarchical feature maps. EviT (Liang et al. 2022) enhances this approach by selecting the top K tokens with the highest average values across all heads for the next stage, merging the remaining tokens. PS-ViT (Yue et al. 2021) further refines the process by iteratively adjusting patch centers towards the object to enrich object information within the hierarchical feature maps. Token Merging (Bolya et al. 2022) employs cosine similarity to progressively merge similar tokens, thereby increasing model throughput.
Fixed-grid methods are constrained by their singular and relatively small receptive fields, and excessively enlarging the grid size leads to increased computational overhead. Dynamic feature-based methods, while adaptive, may discard low-scoring tokens that contain valuable information and often lack end-to-end training capabilities. Our proposed Stepwise Patch Merging approach distinguishes itself from both fixed-grid and dynamic feature-based methods. The Multi-Scale Aggregation module in Stepwise Patch Merging provides an expanded and enriched receptive field, which is advantageous for long-distance modeling. Additionally, the Guided Local Enhancement module enhances the extraction of local discriminative features and supports end-to-end training. Moreover, Stepwise Patch Merging can be directly applied to dense prediction tasks, resulting in improved performance.
Methodology
@@
Multi-Scale Aggregation
@@
Inspired by these mechanisms, the MSA module first divides the input channels $C$ into $N$ distinct heads, each undergoing depth-wise convolutions with varying receptive fields. This method not only reduces the parameter count and computational cost but also facilitates the extraction of multi-granularity information, akin to how different neurons in the brain handle information processing. Subsequently, the MSA module employs larger convolutional kernels to further expand the receptive field, thereby enhancing its capability to model long-range dependencies. Following this, Channel Shuffle (Zhang et al. 2018) is used to interleave channels containing features of different scales, followed by a series of linear projections to fuse these multi-scale features. The number of linear projections is $\frac{C}{N}$, with each projection having unique parameters. Finally, the $N$ heads are concatenated, and a final linear projection adjusts the number of channels to the specified $C^{\prime}$.
@@
Guided Local Enhancement
@@
It is worth noting that when the $[Guide]$ token is stripped of its semantic information, it degrades into the $[CLS]$ token in the vanilla vision transformer (Dosovitskiy et al. 2020). Experiments show that using our $[Guide]$ tokens results in higher performance (see Tab. 6).
Experiments
Image Classification on ImageNet-1K
@@
ResNet18 (He et al. 2016) 31.2 34.0 54.0 36.7 - - - 31.2 51.0 32.7
PVT-Tiny (Wang et al. 2021) 32.9 36.7 59.2 39.3 21.6 39.2 49.0 35.1 56.7 37.3
PVT-Tiny (HAFA) (Chen et al. 2023) 34.5 39.8 62.6 43.3 23.3 42.7 53.3 37.1 59.4 39.3
@@
ResNet50 (He et al. 2016) 44.2 38.0 58.6 41.4 - - - 34.4 55.1 36.7
PVT-Small (Wang et al. 2021) 44.1 40.4 62.9 43.8 22.9 43.0 55.4 37.8 60.1 40.3
PVT-Tiny (HAFA) (Chen et al. 2023) 45.8 41.8 64.4 45.7 26.0 44.6 56.1 38.9 61.5 41.9
@@
ResNet101 (He et al. 2016) 63.2 40.4 61.1 44.2 - - - 36.4 57.7 38.8
ResNeXt101-32$\times$4d (Xie et al. 2017) 62.8 41.9 62.5 45.9 - - - 37.5 59.4 40.2
PVT-Medium (Wang et al. 2021) 63.9 42.0 64.4 45.6 - - - 39.0 61.6 42.1
@@
Setting. We first evaluate the proposed SPM framework on the ImageNet-1K dataset (Deng et al. 2009), which comprises 1.28 million training images and 50,000 validation images spanning 1,000 categories. To ensure a fair comparison, all models are trained on the training set and report the top-1 error rate on the validation set. For data augmentation, we apply a suite of techniques including random cropping, random horizontal flipping (Szegedy et al. 2015), label smoothing regularization (Szegedy et al. 2016), mixup (Zhang et al. 2017), CutMix (Yun et al. 2019), and random erasing (Zhong et al. 2020). These augmentations are employed to enhance the robustness and generalization ability of the models. During training, we use the AdamW optimizer (Loshchilov and Hutter 2017) with a momentum parameter of 0.9, a mini-batch size of 128, and a weight decay of $5\times 10^{-2}$. The initial learning rate is set to $1\times 10^{-3}$ and follows a cosine annealing schedule (Loshchilov and Hutter 2016) to gradually reduce the learning rate. All models are trained from scratch for 300 epochs on eight NVIDIA A100 GPUs. For evaluation, we adopt the standard center crop strategy on the validation set, where a $224\times 224$ patch is extracted from each image to assess the classification accuracy.
Backbone #Params (M) Top-1 Acc. (%)
ResNet18 (He et al. 2016) 11.7 68.5
DeiT-Tiny/16 (Touvron et al. 2021) 5.7 72.2
PVT-Tiny (Wang et al. 2021) 13.2 75.1
PVT-Tiny (HAFA) (Chen et al. 2023) 14.6 77.5
PVT-Tiny (SPM) 14.0 79.5 (+4.4)
ResNet50 (He et al. 2016) 25.6 78.5
ResNeXt50-32$\times$4d (Xie et al. 2017) 25.0 79.5
DeiT-Small/16 (Touvron et al. 2021) 22.1 79.9
HRNet-W32 (Wang et al. 2020) 41.2 78.5
PVT-Small (Wang et al. 2021) 24.5 79.8
PVT-Small (HAFA) (Chen et al. 2023) 25.8 80.1
PVT-Small (SPM) 25.3 81.7 (+1.9)
ResNeXt101-64$\times$4d (Xie et al. 2017) 83.5 81.5
ViT-Base/16 (Dosovitskiy et al. 2020) 86.6 81.8
DeiT-Base/16 (Touvron et al. 2021) 86.6 81.8
PVT-Medium (Wang et al. 2021) 44.2 81.2
PVT-Medium (SPM) 45.0 81.9 (+0.7)
@@
Result. In Tab. 2, we observe that incorporating the SPM framework into the PVT results in significant improvements in classification accuracy, specifically by 4.4%, 1.9%, and 0.7% in the Tiny, Small, and Medium models, respectively, while adding only a minimal number of parameters compared to the original PVT. The final experimental results indicate that the accuracy of models of various sizes has been enhanced, with the most notable improvement observed in the Tiny model. Remarkably, the combination of PVT-Tiny and SPM achieved a top-1 accuracy of 79.5%, which is comparable to the performance of PVT-Small, despite PVT-Small having nearly 70% more parameters than PVT-Tiny. Furthermore, with the integration of SPM, PVT-Small surpassed PVT-Medium by 0.5%.
Object Detection on COCO

### v2

@@ Abstract
The hierarchical architecture has become a mainstream design paradigm for Vision Transformers (ViTs), with Patch Merging serving as the pivotal component that transforms a columnar architecture into a hierarchical one. Drawing inspiration from the brain’s ability to integrate global and local information for comprehensive visual understanding, we propose Stepwise Patch Merging (SPM), which enhances the subsequent attention mechanism’s ability to ’see’ better. SPM consists of Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE) striking a proper balance between long-range dependency modeling and local feature enhancement. Extensive experiments conducted on benchmark datasets, including ImageNet-1K, COCO, and ADE20K, demonstrate that SPM significantly improves the performance of various models, particularly in dense prediction tasks such as object detection and semantic segmentation. Meanwhile, experiments show that combining SPM with different backbones can further improve performance. The code has been released at https://github.com/Yonghao-Yu/StepwisePatchMerging.
1 Introduction
@@
Transformers have demonstrated remarkable advancements in natural language processing (NLP) Vaswani et al. (2017); Devlin et al. (2018), and their application has recently extended significantly into the computer vision (CV) domain Dosovitskiy et al. (2020). To enhance their adaptability to downstream tasks, hierarchical vision transformers (HVTs) Wang et al. (2021); Liu et al. (2021) have been developed. These architectures draw inspiration from the pyramid structure utilized in convolutional neural networks (CNNs) Krizhevsky et al. (2012); He et al. (2016). In HVTs, transformer blocks are segmented into multiple stages, resulting in a progressive reduction of feature map sizes and an increase in the number of channels as the network depth increases.
@@
Figure 1: Overview of the proposed Stepwise Patch Merging (SPM) framework built upon the HVTs. The SPM framework comprises two sequential modules: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE).
@@
HVTs commonly utilize either standard convolutional layers or linear projection layers to amalgamate adjacent tokens Wang et al. (2021); Liu et al. (2021); Ren et al. (2022), with the objective of generating hierarchical feature maps. Nevertheless, fixed-grid methods can limit the representational capacity of vision transformers in modeling geometric transformations, as not every pixel contributes equally to an output unit Luo et al. (2016). To overcome this limitation, adaptive methods have been proposed to derive more informative downsampled tokens for subsequent processing. For example, LIT Pan et al. (2022), drawing inspiration from deformable convolutions Dai et al. (2017), learns a grid of offsets to adaptively adjust spatial sampling locations for merging neighboring patches from a sub-window in a feature map. Similarly, TCformer Zeng et al. (2022) employs a variant of the k-nearest neighbor-based density peaks clustering algorithm (DPC-KNN) Du et al. (2016) to aggregate redundant patches, generating more patches on the target object to capture additional information. HAFA Chen et al. (2023) integrates the methodologies of LIT and TCformer, predicting offsets to adjust the sampling center of patches in the shallow layers, while employing clustering in the deeper layers to group patches with similar semantics in the feature space. However, these methods face several common challenges. They often exhibit limited capacity for modeling long-distance relationships and suffer from a loss of spatial information due to the clustering process. Additionally, the clustering algorithms used are typically not amenable to end-to-end training, leading to inefficiencies. The integration of multiple modules, as seen in HAFA, further complicates their generalizability across different applications.
@@
In the brain, the primary visual cortex (V1), integral to initial visual processing, houses neurons with relatively small receptive fields that are crucial for detecting fine, localized visual features such as edges and orientations. As visual information propagates to higher cortical areas like V2, V3, and V4, neurons with increasingly larger receptive fields integrate these initial perceptions, facilitating the recognition of more complex patterns and broader contextual elements Hubel and Wiesel (1962). Additionally, the visual cortex benefits from a dynamic feedback system where higher-order areas like the inferotemporal cortex (IT) provide contextual modulation to lower areas. This top-down modulation is essential for refining the perception of local features within their broader environmental matrix, enhancing both the accuracy and relevance of visual processing Gilbert and Li (2013).
@@
Inspired by the orchestrated activities across various cortical areas, we introduce Stepwise Patch Merging (SPM), as shown in Fig. 1, a novel approach designed to enhance the receptive field while preserving local details. SPM framework consists of two sequential stages: Multi-Scale Aggregation (MSA) and Guided Local Enhancement (GLE). In the MSA stage, spatial dimensions are preserved while channel dimensions are increased to a designated size. This process aggregates multi-scale information, enriching the semantic content to accommodate the increased capacity of the feature map. Subsequently, the GLE stage reduces the spatial dimensions of the feature map while maintaining the channel dimensions. Given that the input to GLE already contains rich semantic information, this stage emphasizes local information, optimizing it for downstream dense prediction tasks such as object detection and semantic segmentation. The distinct focus and reasonable division of labor between the MSA and GLE modules ensure that the SPM architecture serves as a flexible, drop-in replacement for existing HVTs.
@@
  • We propose an innovative technique termed Stepwise Patch Merging (SPM), which serves as a plug-in replacement within HVTs, leading to substantial performance enhancements.
@@
2 Related Work
2.1 Vision Transformer
The Vision Transformer (ViT) Dosovitskiy et al. (2020) revolutionized visual tasks by introducing the transformer architecture to computer vision. ViT segments images into non-overlapping patches, projects these patches linearly into token sequences, and processes them using a transformer encoder. ViT models have demonstrated superior performance in image classification and other downstream tasks, surpassing CNNs He et al. (2016); Krizhevsky et al. (2012) when trained with large-scale pretraining datasets and advanced training methodologies. Motivated by the success of CNNs and the necessity to address dense prediction tasks, researchers have incorporated the feature pyramid structure within transformers. This innovation has led to the development and widespread adoption of HVTs Dong et al. (2022); Ren et al. (2022); Guo et al. (2022); Hou et al. (2022); Lin et al. (2023).
2.2 Hierarchical Feature Representation
Hierarchical feature representation plays a pivotal role in dense prediction tasks, prompting extensive research in this domain. Existing approaches can be broadly categorized into fixed-grid and dynamic feature-based methods. Fixed-grid methods, exemplified by works such as PVT Wang et al. (2021) and Swin Liu et al. (2021), merge patches within adjacent windows using 2D convolution. In contrast, dynamic methods, such as DynamicViT Rao et al. (2021) adaptively extract features by eliminating redundant patches and retaining essential ones, thereby forming hierarchical feature maps. EviT Liang et al. (2022) enhances this approach by selecting the top K tokens with the highest average values across all heads for the next stage, merging the remaining tokens. PS-ViT Yue et al. (2021) further refines the process by iteratively adjusting patch centers towards the object to enrich object information within the hierarchical feature maps. Token Merging Bolya et al. (2022) employs cosine similarity to progressively merge similar tokens, thereby increasing model throughput.
Fixed-grid methods are constrained by their singular and relatively small receptive fields, and excessively enlarging the grid size leads to increased computational overhead. Dynamic feature-based methods, while adaptive, may discard low-scoring tokens that contain valuable information and often lack end-to-end training capabilities. Our proposed SPM distinguishes itself from both fixed-grid and dynamic feature-based methods.
2.3 Global/Coarse and Local/Fine Feature Fusion
Fusing coarse- and fine-grained features is not novel; however, most prior work treats both feature types equally and performs simple fusion. For example, PLG-ViT Ebert et al. (2023) fuses global and local features using the Hadamard product. SPM uses a lightweight module (GTG) to generate global information that guides the generation of detailed features, an approach demonstrated by experiments to be highly effective.
3 Methodology
@@
3.1 Multi-Scale Aggregation
@@
Inspired by these mechanisms, the MSA module first divides the input channels $C$ into $N$ distinct heads, each undergoing depth-wise convolutions with varying receptive fields. This method not only reduces the parameter count and computational cost but also facilitates the extraction of multi-granularity information, akin to how different neurons in the brain handle information processing. Subsequently, the MSA module employs larger convolutional kernels to further expand the receptive field, thereby enhancing its capability to model long-range dependencies. Following this, Channel Shuffle Zhang et al. (2018) is used to interleave channels containing features of different scales, followed by a series of linear projections to fuse these multi-scale features. The number of linear projections is $\frac{C}{N}$, with each projection having unique parameters. Finally, the $N$ heads are concatenated, and a final linear projection adjusts the number of channels to the specified $C^{\prime}$.
@@
3.2 Guided Local Enhancement
@@
It is worth noting that when the $[Guide]$ token is stripped of its semantic information, it degrades into the $[CLS]$ token in the vanilla vision transformer Dosovitskiy et al. (2020). Experiments show that using our $[Guide]$ tokens results in higher performance (see Tab. 6).
4 Experiments
4.1 Image Classification on ImageNet-1K
Setting.
We first evaluate the proposed SPM framework on the ImageNet-1K dataset Deng et al. (2009), which comprises 1.28 million training images and 50,000 validation images spanning 1,000 categories. To ensure a fair comparison, all models are trained on the training set and report the top-1 error rate on the validation set. For data augmentation, we apply a suite of techniques including random cropping, random horizontal flipping Szegedy et al. (2015), label smoothing regularization Szegedy et al. (2016), mixup Zhang et al. (2017), CutMix Yun et al. (2019), and random erasing Zhong et al. (2020). These augmentations are employed to enhance the robustness and generalization ability of the models. During training, we use the AdamW optimizer Loshchilov and Hutter (2017) with a momentum parameter of 0.9, a mini-batch size of 128, and a weight decay of $5\times 10^{-2}$. The initial learning rate is set to $1\times 10^{-3}$ and follows a cosine annealing schedule Loshchilov and Hutter (2016) to gradually reduce the learning rate. All models are trained from scratch for 300 epochs on eight NVIDIA A100 GPUs. For evaluation, we adopt the standard center crop strategy on the validation set, where a $224\times 224$ patch is extracted from each image to assess the classification accuracy.
Backbone #Params (M) Top-1 Acc. (%)
ResNet18 He et al. (2016) 11.7 68.5
DeiT-Tiny/16 Touvron et al. (2021) 5.7 72.2
PVT-Tiny Wang et al. (2021) 13.2 75.1
PVT-Tiny (HAFA) Chen et al. (2023) 14.6 77.5
PVT-Tiny (SPM) 14.0 79.5 (+4.4)
ResNet50 He et al. (2016) 25.6 78.5
ResNeXt50-32$\times$4d Xie et al. (2017) 25.0 79.5
DeiT-Small/16 Touvron et al. (2021) 22.1 79.9
HRNet-W32 Wang et al. (2020) 41.2 78.5
PVT-Small Wang et al. (2021) 24.5 79.8
PVT-Small (HAFA) Chen et al. (2023) 25.8 80.1
PVT-Small (SPM) 25.3 81.7 (+1.9)
ResNeXt101-64$\times$4d Xie et al. (2017) 83.5 81.5
ViT-Base/16 Dosovitskiy et al. (2020) 86.6 81.8
DeiT-Base/16 Touvron et al. (2021) 86.6 81.8
PVT-Medium Wang et al. (2021) 44.2 81.2
PVT-Medium (SPM) 45.0 81.9 (+0.7)
@@ 4.1 Image Classification on ImageNet-1K
ResNet18 He et al. (2016) 31.2 34.0 54.0 36.7 - - - 31.2 51.0 32.7
PVT-Tiny Wang et al. (2021) 32.9 36.7 59.2 39.3 21.6 39.2 49.0 35.1 56.7 37.3
PVT-Tiny (HAFA) Chen et al. (2023) 34.5 39.8 62.6 43.3 23.3 42.7 53.3 37.1 59.4 39.3
@@
ResNet50 He et al. (2016) 44.2 38.0 58.6 41.4 - - - 34.4 55.1 36.7
PVT-Small Wang et al. (2021) 44.1 40.4 62.9 43.8 22.9 43.0 55.4 37.8 60.1 40.3
PVT-Tiny (HAFA) Chen et al. (2023) 45.8 41.8 64.4 45.7 26.0 44.6 56.1 38.9 61.5 41.9
@@
ResNet101 He et al. (2016) 63.2 40.4 61.1 44.2 - - - 36.4 57.7 38.8
ResNeXt101-32$\times$4d Xie et al. (2017) 62.8 41.9 62.5 45.9 - - - 37.5 59.4 40.2
PVT-Medium Wang et al. (2021) 63.9 42.0 64.4 45.6 - - - 39.0 61.6 42.1
@@
Result.
@@
In Tab. 1, we observe that incorporating the SPM framework into the PVT results in significant improvements in classification accuracy, specifically by 4.4%, 1.9%, and 0.7% in the Tiny, Small, and Medium models, respectively. The final experimental results indicate that the accuracy of models of various sizes has been enhanced, with the most notable improvement observed in the Tiny model. Furthermore, with the integration of SPM, PVT-Small surpassed PVT-Medium by 0.5%.
4.2 Object Detection on COCO
Setting.